OpenStudy (anonymous):

Match the power with its value. 5^2 A. - 1/25 5^-2 B. 1/25 -5^2 C. -25 -5^-2 D. 25

OpenStudy (anonymous):

5^2=25

OpenStudy (anonymous):

-5^2=-25

OpenStudy (anonymous):

5^-2=1/25

OpenStudy (anonymous):

-5^-2=-1/25 thats it.

OpenStudy (anonymous):

TYSM

OpenStudy (anonymous):

ur's welcome

OpenStudy (anonymous):

appreciate me by meday :D

OpenStudy (anonymous):

sorry by medal

OpenStudy (anonymous):

its ok thank you again :)

OpenStudy (anonymous):

thnks dear.

OpenStudy (anonymous):

yw :D

OpenStudy (anonymous):

:)